# Pressure2Motion: Hierarchical Motion Synthesis from Ground Pressure with Text Guidance

### v1

@@ Introduction
Finally, to ensure a strong coupling between the generated motion and the input pressure signal, we introduce a pressure-motion consistency loss alongside the standard diffusion loss. This loss measures the alignment between the synthesized motion and the pressure features. Since pressure primarily reflects foot contact and overall body trajectory, we compute the consistency loss based on the global positions of five key joints: the pelvis root, left/right ankle, and left/right foot. The consistency loss is given by:
@@
Experiments
Experimental Settings
Dataset.
Figure 4: Visual comparisons on the MPL dataset. Yellow denotes the predicted results of different methods; blue represents the ground-truth motions. The motions generated by ours align best with the ground truth, especially in the foot region.
Method FID $\downarrow$ Foot Skating $\downarrow$ LMPJPE $\downarrow$ MPJPE $\downarrow$
@@
Ground Truth 0.002 0.0000 0.0000 0.0000 0.000 0.64
MDM 4.819 0.1029 0.2550 0.2996 0.2744 0.458
MotionDiffuse 3.812 0.1138 0.2305 0.2884 0.2650 0.486
OmniControl 0.315 0.0629 0.1362 0.1719 0.1035 0.523
MaskControl 0.388 0.0617 0.1335 0.1695 0.1009 0.534
Regression 40.015 0.7338 0.4322 0.4896 0.5864 0.127
Ours 0.262 0.0553 0.1273 0.1622 0.1445 0.545
Table 1: Comparison of motion synthesis with pressure control signal on the MPL dataset.
We train and evaluate our approach on the proposed MPL dataset. Data augmentations, including random spatial translations and rotations, are implemented to enhance generalization. The dataset is split into training, validation, and test sets with proportions of 80%, 15%, and 5%. More training details can be found in the Appendix.
@@
We follow the evaluation protocol from OmniControl (omnicontrol), combining motion quality from HumanML3D (t2m) and trajectory accuracy from GMD (karunratanakul2023guided). Additionally, we introduce two alignment metrics—Mean Per Joint Position Error (MPJPE) and Lower-body MPJPE (LMPJPE)—to assess the consistency between the synthesized motion and ground pressure signals, with LMPJPE focusing on joints relevant to foot-ground interactions. All evaluations are conducted using a motion evaluator trained on our MPL dataset, following HumanML3D settings.
Comparison Experiments
As no prior work has explored human motion synthesis conditioned on pressure sequences, we reimplement several state-of-the-art motion generation models as baselines, including diffusion-based models (tevet2022human; motiondiffuse; omnicontrol) and masked generative models (maskcontrol). To ensure fair comparisons and enable per-frame pressure-based control, we adapt each baseline accordingly. For MDM and MotionDiffuse, we concatenate the global and local pressure embeddings and append them to the noisy motion input at each denoising step. This allows these models to incorporate pressure signals at each frame, providing a consistent pressure-aware motion synthesis. For OmniControl and MaskControl, we replace the original spatial control inputs with the combined pressure embeddings, enabling these models to condition on pressure in a comparable manner to our approach.
In addition to these baselines, we modify the diffusion process in our method by reducing the multi-step denoising process to a single-step denoising. This adaptation transforms our method into a discriminative regression model. By shifting to a regression-based approach, we aim to investigate the effectiveness of a generative model (which is our approach) in synthesizing motion from sparse pressure data.
The results in Table 1 demonstrate that our method outperforms all baselines across various metrics. We achieve the lowest FID (0.262), indicating our generated motion sequences are closer to real-world data. Our model also significantly reduces foot skating (0.0553), and leads in both MPJPE (0.1622) and LMPJPE (0.1273), showcasing better spatial accuracy, particularly in foot-ground interactions. In terms of R-precision, we outperform baselines with a score of 0.545, demonstrating better motion retrieval performance.
As visualized in Figure 4, our method produces physically realistic motions with accurate foot-ground contact, avoiding the foot-sliding and joint-misalignment artifacts common in baselines.
@@
TrajErr
(>>50cm) $\downarrow$
w/o MT 0.543 0.0665 0.1943 0.2357 0.3086
w/o PS 0.847 0.0629 0.1555 0.2025 0.2041
w/o CL 0.282 0.0721 0.1550 0.1896 0.1621
Full 0.262 0.0553 0.1273 0.1622 0.1445
Table 2: Ablation study of: Movement Trajectory (MT), Posture Shifts (PS), and the Consistency Loss (CL).
Figure 5: Visualization results of ablation study.
Ablation Study
Our ablation study evaluates each key component, with quantitative results in Table 2 and qualitative examples in Figure 5. Omitting the Pressure-Inferred Movement Trajectory causes significant degradation in alignment-related metrics. Removing Posture Shifts significantly degrades realism (FID), which demonstrates the need for fine-grained adjustments. Furthermore, removing the Consistency Loss increases foot skating, confirming its role in ensuring plausible foot-ground contact. Our complete method outperforms all ablated versions, demonstrating that each component is crucial for synthesizing high-fidelity motion.
Effect of Text on Motion Synthesis
@@
We conducted experiments to evaluate how varying textual inputs, alongside a pressure-only input condition (devoid of any textual content), influence the outcomes of motion synthesis. Figure 6 illustrates the visualization results, where (a) represents the motion generated with only pressure data, while (b), (c), and (d) correspond to motions generated with different textual descriptions.
@@
In all four cases, the model successfully learns from the pressure signals, demonstrating the model’s ability to incorporate the physical dynamics encoded in the pressure data. Additionally, variations in the textual input result in different motion patterns. Specifically, the influence of text is more pronounced on the upper body, while the motion of the lower body remains closely tied to the pressure data. These findings validate the model’s ability to synthesize distinct motions based on varying textual descriptions while still maintaining alignment with the pressure signals.
@@
Figure 6: Comparison of motion synthesis results: (a) generated from pressure only, versus (b-d) generated from pressure with varying text prompts.
Conclusion
@@
In this work, we introduce Pressure2Motion, a pioneering approach for human motion synthesis from ground pressure and text prompts. Eliminating the need for cameras or wearable devices, our method enables privacy-preserving and non-intrusive motion capture. We tackle the ill-posed nature of this task with a hierarchical diffusion model specifically designed to interpret physical signals; its dual-level feature extractor decodes pressure into both broad movement trajectories and fine-grained posture adjustments. Our experiments establish strong baseline performance, demonstrating that Pressure2Motion synthesizes high-fidelity and plausible motion sequences. Furthermore, our introduction of the MPL dataset provides the first benchmark to spur future research in this new direction. Future work may focus on enhancing robustness and expanding the dataset to include more dynamic activities, while also leveraging pressure data in simulated reinforcement learning and exploring its potential for body shape estimation in personalized motion modeling.
Limitations.
Our current work is constrained by the dataset’s limited diversity of motion types. More complex scenarios involving inclined surfaces or real-world interactions have yet to be explored.
Pressure2Motion:
Hierarchical Motion Synthesis from Ground Pressure with Text Guidance
@@ 1 MPL Dataset Details
The MPL dataset is developed to facilitate research in generating full-body human motion from the highly sparse inputs of ground pressure and text prompts, building on top of the MotionPRO dataset (ren2025motionpro), which contains a large-scale collection of human motion sequences captured using plantar pressure sensors. For our research, we extend this dataset by incorporating textual descriptions.
@@ 1.1 Data Distribution
Figure 7 illustrates the distribution for gender, height, and weight. This coverage enhances the robustness and applicability of our motion synthesis model across real-world variations in body structure.
@@ 1.2 Data Processing
We reformatted the SMPL(SMPL) parameters from the MotionPro dataset into a more comprehensive motion representation following the HumanML3D (t2m) convention. Each motion sequence of length $N$ is transformed into a representation of shape $(N,263)$ where each frame encodes the pelvis velocity, local joint positions, joint velocities, joint rotations (in pelvis space), and binary foot contact indicators.
@@ 1.3 Caption Process
Text descriptions are an integral part of the dataset, providing semantic guidance for the motion generation process. To ensure diversity and semantic richness in textual prompts, descriptions are generated using Qwen2.5-VL (Qwen2-VL), a vision-language model capable of processing long video sequences. Specifically, given a motion clip and a brief action keyword from the original MotionPRO dataset, we provide the RGB video frames and keyword as input to Qwen2.5-VL. The model interprets the human activity within the video context and generates five diverse captions at varying levels of detail. These descriptions range from simple high-level actions (e.g., “The person is walking”) to more intricate and detailed descriptions (e.g., “The person is walking with a slight leftward tilt and right arm movement”).
@@ 2 Implementation Details
Our models are implemented in PyTorch and trained on 8 NVIDIA A800 GPUs for a total of 100,000 iterations. We adopt the AdamW optimizer (loshchilov2017decoupled) with a learning rate of $1\times 10^{-5}$. The ControlNet is initialized with pretrained weights from MDM (tevet2022human). During training, the parameters of the Movement Trajectory extraction module and the pretrained MDM are frozen to retain their original representations.
@@
We follow prior works and use CLIP (radford2021learning) to encode text prompts into 512-dimensional embeddings. The output features of both the ControlNet and the Adapter modules are also of size 512 to ensure compatibility with the pretrained MDM architecture. The Pressure-Inferred Movement Trajectory and Pressure-Induced Posture Shifts are extracted with output dimensions of $(B,L,39)$ and $(B,L,256)$ respectively, where $B$ is the batch size and $L=196$ is the sequence length. The 39-dimensional trajectory representation includes the global 3D positions ($XYZ$) of the root, left/right ankles, and left/right toes (total 5 joints × 3 = 15), as well as 6D rotation representations for the left/right ankles and toes (4 joints × 6 = 24).
@@
To improve robustness to text variations, we randomly drop 10% of the text conditions during training. This enables the use of Classifier-Free Guidance (CFG) (ho2022classifier) during inference, where we apply a CFG scale of 5. We adopt a standard DDPM (ho2020denoising) framework with $T=1000$ denoising steps. The control strength $\tau$ for injecting pressure signals is defined as $\tau=\frac{20\hat{\Sigma}_{t}}{L}$, where $\hat{\Sigma}_{t}=\min(\Sigma_{t},0.01)$. We set $\lambda_{\text{diff}}=1$ and $\lambda_{\text{cons}}=5$ throughout all experiments.
3 More Visualization Results
We present more motion synthesis results to further demonstrate the effectiveness of our method in generating human motion from sparse pressure data. Figure 8 showcases a variety of synthesized motions across different scenarios, highlighting the robustness of our approach in the pressure-to-motion task. The first few cases demonstrate typical human actions, such as walking, standing, and some daily activities. Our model generates high-fidelity motions that align well with the pressure data and text prompts, maintaining both physical consistency and semantic plausibility.
Particularly interesting are the last two cases in Figure 8. The second-to-last case corresponds to a jumping motion, where no pressure is applied during the jump. Despite the absence of pressure in the air, our model successfully generates a realistic jumping motion, demonstrating its ability to handle scenarios with no foot-ground contact.
The final case in Figure 8 shows a dynamic plank position, which involves complex pressure distributions from both the hands and feet. Our method effectively handles this rare and specialized pressure contact scenario, generating a physically plausible motion that corresponds to the simultaneous pressure from all four limbs. This demonstrates the versatility of our model in handling uncommon or intricate pressure patterns.
@@ 4 Evaluation Details
[…]
TrajErr
(>>50cm) $\downarrow$
w/o ctrlnet 1.3683 0.0621 0.1694 0.1951 0.1650
w/o Adapter 0.695 0.0634 0.1702 0.2092 0.1742
Full 0.262 0.0553 0.1273 0.1622 0.1445
Table 3: Additional Ablation study of Model Archtecture.
@@
We evaluate motion quality, physical consistency, and semantic alignment of the generated motions using the following metrics:
  • Fréchet Inception Distance (FID) $\downarrow$: FID measures the distributional distance between generated motions and ground-truth motions in the feature space. In our setting, motions are encoded via a pre-trained motion encoder, and FID is computed on the extracted features. Lower FID indicates that the generated motions are more realistic and distributionally similar to real data.
@@
    where $T$ is the number of frames, $N$ is the number of joints, and $\hat{\mathbf{p}}_{t,j}$ and $\mathbf{p}_{t,j}$ denote the predicted and ground-truth positions of joint $j$ at time $t$. This metric evaluates the spatial alignment between generated and real motions.
  • Lower-body MPJPE (L-MPJPE) $\downarrow$: A variant of MPJPE that only considers lower-body joints (e.g., hips, knees, ankles, feet), which are most relevant to pressure-ground interactions. It reflects the model’s ability to generate physically grounded lower-body motion. Lower is better.
@@
  • R-Precision (Top-3) $\uparrow$: R-Precision measures the semantic consistency between generated motions and their associated text prompts. We use a joint motion-text encoder to compute the similarity between generated motion features and ground-truth text embeddings. R-Precision@3 reflects whether the correct caption is ranked among the top-3 retrieved results for a generated motion. Higher values indicate better semantic alignment.

### v2

@@ 4.3 Hierarchical Pressure-Modulated Motion Synthesis
Finally, to ensure a strong coupling between the generated motion and the input pressure signal, we introduce a pressure-motion consistency loss alongside the standard diffusion loss. This loss measures the alignment between the reconstructed motion’s key joints $R(\hat{\mathbf{x}}_{0}^{\prime})$and the pressure-inferred trajectories $\mathbf{T}_{\text{traj}}$. Since pressure primarily reflects foot contact and overall body trajectory, we compute the consistency loss based on the global positions of five key joints: the pelvis root, left/right ankle, and left/right foot. The consistency loss is given by:
@@
5 Experiments
5.1 Experimental Settings
Figure 4: Visual comparisons on the MPL dataset. Yellow denotes the predicted results of different methods; blue represents the ground-truth motions. The motions reconstructed by ours align best with the ground truth, especially in the foot region.
Table 1: Comparison of motion reconstruction with pressure control signal on the MPL dataset.
Method FID $\downarrow$ Foot Skating $\downarrow$ CoP Error $\downarrow$ LMPJPE $\downarrow$ MPJPE $\downarrow$
@@ 5.1 Experimental Settings
Ground Truth 0.002 0.0000 0.0000 0.0000 0.0000 0.000 0.64
MDM[43] 4.819 0.1029 0.9238 0.2550 0.2996 0.2744 0.458
MotionDiffuse[55] 3.812 0.1138 0.8765 0.2305 0.2884 0.2650 0.486
OmniControl[49] 0.315 0.0629 0.5862 0.1362 0.1719 0.1035 0.523
MaskControl[36] 0.388 0.0617 0.5644 0.1335 0.1695 0.1009 0.534
Text-Only 0.872 0.1560 1.0810 0.2320 0.2838 0.3082 0.2866
Regression 40.015 0.7338 1.4832 0.4322 0.4896 0.5864 0.127
Ours 0.262 0.0553 0.4260 0.1273 0.1622 0.1445 0.545
Dataset.
We train and evaluate our approach on the proposed MPL dataset. Data augmentations, including random spatial translations and rotations, are implemented to enhance generalization. The dataset is split into training, validation, and test sets with proportions of 80%, 15%, and 5%. More training details can be found in the Sup. Mat.
@@
We follow the evaluation protocol from OmniControl [49], combining motion quality from HumanML3D [15] and trajectory accuracy from GMD [24]. As our task is a MoCap problem, we introduce two alignment metrics—Mean Per Joint Position Error (MPJPE) and Lower-body MPJPE (LMPJPE)—to assess the reconstruction accuracy , with LMPJPE focusing on joints relevant to foot-ground interactions. Additionally, to directly validate the pressure-motion consistency—a core challenge not addressed by the metrics above—we introduce the Center of Pressure Error (CoP Error). This metric measures the mean L2 distance between the Center of Pressure calculated from the input pressure map and the Center of Pressure inferred from the reconstructed lower-body joint positions. A lower CoP Error indicates superior physical alignment with the input pressure signal. All evaluations are conducted using a motion evaluator trained on our MPL dataset, following HumanML3D settings.
5.2 Comparison Experiments
As stated in Sec. 4.1, this ill-posed pressure-to-motion reconstruction task is best solved using a generative prior. To ensure fair comparisons and enable per-frame pressure-based control, we adapt each baseline accordingly. We emphasize that our task is fundamentally different from controllable generation [49, 56]. Those methods condition on kinematic abstractions (e.g., keypoints or trajectories) with a direct geometric mapping to the output. In contrast, our method must interpret a physical signal (pressure) that is sparse, noisy, and has no direct kinematic correspondence to the full-body pose, presenting a unique ill-posed reconstruction challenge. The specific adaptation details for each baseline are provided in the Sup. Mat.
We compare our method against several baselines [43, 55, 49, 36], a crucial Text-Only baseline (our model with the pressure branch masked), and a simple Regression model (our diffusion model reduced to a single step). All baselines were adapted for per-frame pressure control; details are in the Sup. Mat.
The results in Tab. 1 strongly validate our approach. Our full model achieves state-of-the-art results across reconstruction accuracy (MPJPE/LMPJPE) and realism (FID/Foot Skating). Crucially, our method demonstrates a substantial improvement in the new pressure-motion consistency metric (CoP Error), proving our hierarchical model uniquely learns to align the reconstructed motion with the physical input signal. Furthermore, the Text-Only baseline confirms the necessity of pressure, as its physical realism metrics (CoP Error, Foot Skating) collapse. Our full model also achieves the highest semantic alignment (R-precision), demonstrating a superior ability to balance both physical and textual constraints.
As visualized in Fig. 4, our method produces physically realistic motions with accurate foot-ground contact, avoiding the foot-sliding and joint-misalignment artifacts common in baselines. An analysis of intricate cases, such as motions involving no foot-ground contact (e.g., jumping) or exhibiting uncommon pressure patterns is provided in the Sup. Mat.
Figure 5: Visualization results of ablation study.
Table 2: Ablation study of: Movement Trajectory (MT), Posture Shifts (PS), and the Consistency Loss (CL), and our Hierarchical design (Hi).
@@ 5.2 Comparison Experiments
w/o MT 0.543 0.0665 0.8840 0.1943 0.2357
w/o PS 0.847 0.0629 0.5864 0.1555 0.2025
w/o CL 0.282 0.0721 0.5320 0.1550 0.1896
w/o Hi 0.345 0.0615 0.5610 0.1311 0.1692
Full 0.262 0.0553 0.4260 0.1273 0.1622
5.3 Ablation Study
Our ablation study evaluates each key component, with quantitative results in Tab. 2 and qualitative examples in Fig. 5. As shown in Tab. 2, removing MT, PS, or CL all lead to significant performance degradation. Finally, to validate our hierarchical design, we add a w/o Hi (Hierarchical) baseline. In this variant, we concatenate the $\mathbf{T}_{\text{traj}}$ and $\mathbf{S}_{\text{shift}}$ features into a single unified representation and feed them into the model through a single (non-hierarchical) branch. This unified model performs significantly worse than our full, hierarchical approach. This confirms that separating and hierarchically injecting the high-level trajectory (MT) and low-level posture (PS) signals is crucial for accurate reconstruction. Our complete method outperforms all ablated versions, demonstrating that each component is crucial for reconstructing high-fidelity motion.
5.4 Effect of Text on Motion Reconstruction
@@
We conducted experiments to evaluate how varying textual inputs, alongside a pressure-only input condition (devoid of any textual content), influence the outcomes of motion reconstruction. Fig. 6 illustrates the visualization results, where (a) represents the motion synthesized with only pressure data, while (b), (c), and (d) correspond to motions synthesized with different textual descriptions.
@@
Specifically, the influence of text is more pronounced on the upper body, while the motion of the lower body remains closely tied to the pressure data. These findings validate the model’s ability to reconstruct distinct motions based on varying textual descriptions while still maintaining alignment with the pressure signals.
@@
Figure 6: Comparison of motion reconstruction results: (a) synthesized from pressure only, versus (b-d) synthesized from pressure with varying text prompts.
5.5 Real-World and OOD Generalization
A critical advantage of our Pressure2Motion is its robustness and applicability in real-world, out-of-distribution (OOD) scenarios. As illustrated in Fig. 7, we demonstrate the model’s successful generalization in an uncontrolled environment, such as a corridor. This deployment validates that our approach maintains high-fidelity and physical realism, highlighting its potential for non-visual motion sensing in residential, clinical, or public spaces.
Figure 7: Real-world deployment.
6 Conclusion
@@
In this work, we introduce Pressure2Motion, a pioneering approach for human motion reconstruction from ground pressure and text prompts. Eliminating the need for cameras or wearable devices, our method enables privacy-preserving and non-intrusive motion capture. We tackle the ill-posed nature of this task with a hierarchical diffusion model specifically designed to interpret physical signals; its dual-level feature extractor decodes pressure into both broad movement trajectories and fine-grained posture adjustments. Our experiments establish strong baseline performance, demonstrating that Pressure2Motion reconstructs high-fidelity and plausible motion sequences. Furthermore, our introduction of the MPL dataset provides the first benchmark to spur future research in this new direction.
Limitations. Our primary limitations are twofold: first, the dataset is constrained to flat-surface motions, lacking complex scenarios like inclined surfaces; and second, the inference and training of out model remain computationally demanding.
@@ 7 MPL Dataset Details
The MPL dataset is developed to facilitate research in reconstructing full-body human motion from the highly sparse inputs of ground pressure and text prompts, building on top of the MotionPRO dataset [40], which contains a large-scale collection of human motion sequences captured using plantar pressure sensors. For our research, we extend this dataset by incorporating textual descriptions.
@@ 7.1 Data Distribution
Figure 8 illustrates the distribution for gender, height, and weight. This coverage enhances the robustness and applicability of our motion reconstruction model across real-world variations in body structure.
@@ 7.2 Data Processing
We reformatted the SMPL[28] parameters from the MotionPro dataset into a more comprehensive motion representation following the HumanML3D [15] convention. Each motion sequence of length $N$ is transformed into a representation of shape $(N,263)$ where each frame encodes the pelvis velocity, local joint positions, joint velocities, joint rotations (in pelvis space), and binary foot contact indicators.
@@ 7.3 Caption Process
Text descriptions are an integral part of the dataset, providing semantic guidance for the motion synthesis process. To ensure diversity and semantic richness in textual prompts, descriptions are generated using Qwen2.5-VL [47], a vision-language model capable of processing long video sequences. Specifically, given a motion clip and a brief action keyword from the original MotionPRO dataset, we provide the RGB video frames and keyword as input to Qwen2.5-VL. The model interprets the human activity within the video context and generates five diverse captions at varying levels of detail. These descriptions range from simple high-level actions (e.g., “The person is walking”) to more intricate and detailed descriptions (e.g., “The person is walking with a slight leftward tilt and right arm movement”).
@@ 8.1 Training Setup
Our models are implemented in PyTorch and trained on 8 NVIDIA A800 GPUs for a total of 100,000 iterations. We adopt the AdamW optimizer [29] with a learning rate of $1\times 10^{-5}$. The ControlNet is initialized with pretrained weights from MDM [43]. During training, the parameters of the Movement Trajectory extraction module $\mathcal{F}_{\text{traj}}$ and the pretrained MDM backbone $\mathcal{F}_{\theta}$ are frozen to retain their original representations.
8.2 Network and Feature Dimensions
@@
We follow prior works and use CLIP [37] to encode text prompts into 512-dimensional embeddings. The output features of both the ControlNet and the Adapter modules are also of size 512 to ensure compatibility with the pretrained MDM architecture.
The Pressure-Inferred Movement Trajectory $\mathbf{T}_{\text{traj}}$ and Pressure-Induced Posture Shifts $\mathbf{S}_{\text{shift}}$ are extracted with output dimensions of $(B,L,39)$ and $(B,L,256)$ respectively, where $B$ is the batch size and $L=196$ is the sequence length. The 39-dimensional trajectory representation includes the global 3D positions ($XYZ$) of the root, left/right ankles, and left/right toes (total 5 joints × 3 = 15), as well as 6D rotation representations for the left/right ankles and toes (4 joints $\times$ 6 = 24).
8.3 Diffusion and Loss Hyperparameters
@@
To improve robustness to text variations, we randomly drop 10% of the text conditions during training. This enables the use of Classifier-Free Guidance (CFG) [18] during inference, where we apply a CFG scale of 5. We adopt a standard DDPM [19] framework with $T=1000$ denoising steps. The control strength $\tau$ for injecting pressure signals is defined as $\tau=\frac{20\hat{\Sigma}_{t}}{L}$, where $\hat{\Sigma}_{t}=\min(\Sigma_{t},0.01)$. We set $\lambda_{\text{diff}}=1$ and $\lambda_{\text{cons}}=5$ throughout all experiments.
8.4 Baseline Adaptations
For MDM[43] and MotionDiffuse[55], we concatenate the global and local pressure embeddings and append them to the noisy motion input at each denoising step. This allows these models to incorporate pressure signals at each frame, providing a consistent pressure-aware motion reconstruction. For OmniControl[49] and MaskControl[36], we replace the original spatial control inputs with the combined pressure embeddings, enabling these models to condition on pressure in a comparable manner to our approach.
8.5 Pressure Feature Extractor Details
The Pressure-Inferred Trajectory $\mathcal{F}_{\text{traj}}$ is essential for capturing the overall movement path and body alignment. To extract this information, we utilize a feature extraction module following the architecture from MotionPro [40], which includes a ResNet-based [17] pressure encoder, a temporal information processor, and a fully connected projection layer. Given the sparsity of pressure maps, where valid values are limited and primarily found under the feet during standing, the pressure encoder utilizes a compact ResNet architecture with small convolutional kernels to focus on the localized pressure regions, despite the large size of the pressure map.  Temporal dynamics are captured using the temporal information processor, which combines a GRU [3] to model long-term dependencies with a self-attention mechanism to capture short-term correlations in the pressure sequence.
@@ 9 Evaluation Details
We evaluate motion quality, motion-pressure consistency, and semantic alignment of the reconstructed motions using the following metrics:
  • Center of Pressure (CoP) Error $\downarrow$: This metric directly measures the physical consistency between the input pressure signal and the reconstructed motion. It is calculated as the mean L2 distance between two CoP time-series:
    Pressure CoP ($CoP_{Pressure}$): Calculated from the input pressure map $P_{n}\in\mathbb{R}^{H\times W}$ at frame $n$. The pixel-space CoP (geometric center) is computed as a weighted average:
    $CoP_{P,x}^{(n)}=\frac{\sum_{i,j}P_{n}(i,j)\cdot j}{\sum_{i,j}P_{n}(i,j)},$
    $CoP_{P,z}^{(n)}=\frac{\sum_{i,j}P_{n}(i,j)\cdot i}{\sum_{i,j}P_{n}(i,j)}.$
    This pixel-space CoP is then transformed into motion-space using a pre-calibrated scale $S$ and offset $O$:
    $CoP_{Pressure}^{(n)}=[CoP_{P,x}^{(n)},0,CoP_{P,z}^{(n)}]\odot S+O.$
    Motion CoP ($CoP_{Motion}$): Inferred from the reconstructed motion’s lower-body joints. We use a softmax-weighted average of the $K$ key foot joints’ (e.g., ankles, toes) ground projections $j_{k}=[j_{k,x},j_{k,y},j_{k,z}]$, where the weight $w_{k}$ is inversely related to the joint’s height $j_{k,y}$:
    $w_{k}^{(n)}=\frac{\exp(-j_{k,y}^{(n)}/\tau)}{\sum_{l=1}^{K}\exp(-j_{l,y}^{(n)}/\tau)},$
    $CoP_{Motion}^{(n)}=\sum_{k=1}^{K}w_{k}^{(n)}\cdot[j_{k,x}^{(n)},0,j_{k,z}^{(n)}].$
    CoP Error ($\mathcal{L}_{CoP}$): The final error is the mean Euclidean distance over all $N$ frames and $B$ batch items:
    $\mathcal{L}_{CoP}=\frac{1}{B\cdot N}\sum_{b=1}^{B}\sum_{n=1}^{N}\left\|CoP_{Pressure}^{(b,n)}-CoP_{Motion}^{(b,n)}\right\|_{2}.$
    A lower value indicates superior motion-pressure consistency.
  • Fréchet Inception Distance (FID) $\downarrow$: FID measures the distributional distance between reconstructed motions and ground-truth motions in the feature space. In our setting, motions are encoded via a pre-trained motion encoder, and FID is computed on the extracted features. Lower FID indicates that the reconstructed motions are more realistic and distributionally similar to real data.
@@
    where $T$ is the number of frames, $N$ is the number of joints, and $\hat{\mathbf{p}}_{t,j}$ and $\mathbf{p}_{t,j}$ denote the predicted and ground-truth positions of joint $j$ at time $t$. This metric evaluates the spatial alignment between reconstructed and real motions.
  • Lower-body MPJPE (L-MPJPE) $\downarrow$: A variant of MPJPE that only considers lower-body joints (e.g., hips, knees, ankles, feet), which are most relevant to pressure-ground interactions. It reflects the model’s ability to reconstruct physically grounded lower-body motion. Lower is better.
@@
  • R-Precision (Top-3) $\uparrow$: R-Precision measures the semantic consistency between reconstructed motions and their associated text prompts. We use a joint motion-text encoder to compute the similarity between reconstructed motion features and ground-truth text embeddings. R-Precision@3 reflects whether the correct caption is ranked among the top-3 retrieved results for a reconstructed motion. Higher values indicate better semantic alignment.